You are developing a procedure for determining traces of lead in paint chips using an acid digestion followed by measurement by atomic absorption spectrophotometry. In order to test the validity of the method, you obtain a NIST (National Institute of Standards and Technology) standard reference material and analyze this material. Five replicas are sampled and analyzed, and the mean of the results is found to be 10.8 ppm with a standard deviation of 0.7 ppm. The value for the standard reference material is know to be 11.7 ppm. Does your method give a statistically correct value at the 95% confidence level?
Answer #1

Atomic absorption spectrophotometry - specially Inductive coupled plasma atomic absorption spectrophotometry (ICPAS) are used to detect heavy metal ions.

Now,

(10.8/11.7) x 100 = 92.3%

but, standard deviation is 0.7,

so, 10.8+0.7= 11.5

(11.5/11.7) x 100 = 98.3%

Yes, this method give a statistically correct value at the 95% confidence level.